# Language-Specific Latent Process Hinders Cross-Lingual Performance

### v2

@@ 4.2 Multilingual processing in the latent space
(b) Mean accuracy in the latent space across languages. Correct answers can be partially identified in the hidden states. Chance-level accuracy is marked with a dashed line.
@@
To understand cross-lingual consistency and knowledge transfer, our next analyses explore the latent multilingual processing by computing the normalized probabilities of answer choices in their native form (nat.) and their English equivalents (en); based on Equation 7.
The log ratios between probability of the native and English answer choices in Figure 1(a) reveal that larger models tend to assign more mass to answers in the native form in the intermediate layers, as $\log p(\textrm{nat. choices})/p(\textrm{en choices})>0$. At layer 22 and beyond, the probability of the native language exceed that of English in Gemma 27B. We see a similar trend in Qwen 7B and 32B with English at layers 4 and 40 respectively. This indicates that large-scale language models operate in individual space that is closer to the token embeddings of the native language. Our results also suggest the contrary for smaller models: Gemma 2B/9B and Qwen 3B. Early layers of the models are similar to the embeddings of the native languages. The hidden states then shift to working predominantly in the English in the middle layers, where the ratio is below zero, before moving closer to the native language space again at the highest layers. This is in line with previous findings [29, 41] that most language-specific neurons are developed at the top and bottom layers of LLMs.
We further interpret the implicit steps taken across layers, by computing for all languages the accuracy of $x^{\ell}_{\textrm{en}}$ and $x^{\ell}_{x}$. In Figure 1(b), the latent predictions are shown to be noisy, however, exhibit accuracy above chance level in higher layers across data sets. Nonetheless, English latent predictions appear more accurate and stable in most cases.
[…]
Belebele mMMLU mHellaSwag
[…]
Gemma 2 2B $-0.6$∗∗∗ $-0.7$∗∗∗ $-0.8$∗∗∗ $-0.7$ $-0.9$∗∗∗ $-0.9$∗∗∗ $-0.2$ $-1.0$∗∗∗ $-1.0$∗∗∗
Gemma 2 9B $-0.5$∗∗∗ $-0.6$∗∗∗ $-0.6$∗∗∗ $-0.3$ $-0.5$ $-0.5$ $0.2$ $0.0$ $0.1$
Gemma 2 27B $-0.3$∗∗∗ $-0.1$ $-0.0$ $-0.6$ $-0.4$ $-0.4$ $-0.8$ $-0.9$∗∗ $-0.9$∗∗
[…]
Qwen 2.5 7B $-0.2$ $-0.2$ $-0.1$ $-0.7$∗∗ $-0.6$ $-0.6$ $-0.9$∗∗∗ $-0.9$∗∗∗ $-1.0$∗∗∗
Qwen 2.5 32B $-0.2$ $0.0$ $-0.0$ $-0.5$ $-0.4$ $-0.5$ $-0.9$∗∗ $-0.9$∗∗ $-0.9$∗∗
Taken together, our analyses imply that smaller models are limited to working within a shared representation, that is closest to English. Conversely, larger models transform the hidden states into subspace of individual languages, rather than the dominant representation. While it is possible for models to reason in a particular language, the process also hinders the models’ ability to infer from the other languages, contributing to cross-lingual inconsistency.
@@ 4.3 Modulating cross-lingual consistency with activation steering
(a) PCA projections of layer activations in Gemma 2 9B.
(b) PCA projections of layer activations in Gemma 2 27B.
We have inferred in Section 4.2 that larger models process multilingual queries differently than smaller models. Figure 3 provides a visual comparison of the hidden activations of 50 parallel Belebele questions between Gemma 9B and 27B, where the dimension of the vectors is reduced with the projection of Principal Component Analysis (PCA) [33]. [^2] PCA is useful to assess if the hidden representations are linearly separable, which predetermines the efficacy of steering vectors [23]. While the activations can be grouped by languages in the early layers in Gemma 9B (Figure 2(a)), they are mostly aligned between layers 18 and 38, and are separated by their written scripts (Latin vs. non-Latin) in the final layer. We note however that activations of low-resource languages, such as Yoruba (blue squares), are the last to be mapped and continue to be disjoint from the other languages across hidden layers. [^3] We include analogous comparison between Qwen 3B and 32B in Appendix C.1.
Figure 4: Steered Gemma 2B. Note that $\gamma>0$ indicates an activation vector that pushes the model towards English processing in the latent space, and vice versa.
(a) Steered models ($\gamma=2.0$) evaluated on Belebele.
(b) Steered models ($\gamma=2.0$) evaluated on mMMLU.
We extract steering vectors from 50 randomly sampled parallel questions from each data set, and evaluate the cross-lingual performance of the steered models on held-out data. Figure 4 reports a sweep of the steering vector multiplier, $\gamma$, on layers 6 and 10 of Gemma 2B. Positively steered models become increasingly accurate (peaking at $2\leq\gamma\leq 3$), are more consistent with English and receive more positive transfer from English, as the hidden states move towards English subspace. [^4] We observe generally consistent variability across $\gamma$, in terms of standard deviations and confidence intervals over languages. This is omitted in the graph to avoid visual clutter. Similar plots for Qwen models are included in Appendix C.2.
We also observe that steering vectors at the middle layers are the most impactful, because the hidden representations are intended to be aligned at this stage (Figure 3), to allow for language-agnostic task solving [41]. [^5] The performance of steered Gemma models across layers is reported in Figure 8 in Appendix C.3. Figure 5 shows the accuracy changes (%) in steered models across languages. In Figure 4(a), we see low-resource (accuracy) languages written in Latin script benefit the most from the method in Belebele; whereas the improvement in mMMLU (Figure 4(b)) is less notable without steering vectors specialized to each topics. [^6] Similar result can be observed in mHellaSwag, which is reported in Figure 9 in the Appendix. We also include corresponding analyses on positive transfer and consistency in Appendix C.4. The method is also less effective on larger models. We speculate on two possible reasons: i) the performance of larger models is saturated, with little headroom for improvement; ii) the hidden representations might become increasingly non-linear with scale, therefore linear perturbations with steering vectors are less useful on larger models.
@@ 5 Related work
Qi et al. [20] define pairwise cross-lingual knowledge consistency as the equality in ranked outputs scaled by softmax, whereas Ifergan et al. [12] and Goldman et al. [7] measure the extent to which a language pair are accurate in a similar way with factual questions. Our use of Spearman’s correlation in the evaluation of models’ consistency allows a more direct approach, while also taking into account the relative order of ranks. Across languages, script similarity has been proposed to be a major factor to drive knowledge transfer [20, 12, 7]. Our work also supports previous findings that models are capable of processing and storing information with independent representation of different languages [12], and explains previous findings that transfer performance increases with model size only up to 14B for Qwen 2.5, indicating the ability of larger models to retrieve knowledge learned from independent languages [7].
@@
A growing body of work has shown that multilingual LLMs use English as a pivot language in the hidden layers [32], and share the subspace across languages and modalities [34]. Alabi et al. [1] find that the phenomenon also applies to models adapted to new languages, where finetuned adapters preserve most of the representation of the source language of the model, rather than working in a parallel isolated structure. Other works suggest the distinction between language-agnostic and language-specific neurons [41, 31], motivating a series of methods that improve multilingual representation and performance of LLMs with targeted supervision [40, 11]. The current study extends this understanding of multilingual models, and provides evidence for larger LLM’s capability to encode meaningful representation in languages beyond English.
@@
Prior work has shown that explicit alignment with English reasoning improves multilingual performance. Existing approaches include English translations of multilingual queries [27, 43, 16], using English reasoning traces [27, 21, 37, 16], post-hoc translation from English answers [16], and trained alignment with English representation [38, 10, 26]. Our exploration on cross-lingual activation steering adds to this line of work, and offers a straightforward method to tap into English reasoning capability without training or additional inference costs.
@@ 6 Discussion
This work investigates the challenges of cross-lingual transfer and consistency in LLMs. We examine the implicit processing of multilingual queries and show LLMs can operate in a multilingual subspace. With more capacity, larger models exhibit a higher utilization of independent subspaces of individual languages, whereas smaller models are constrained to work predominantly in the shared semantic space. Models are also less likely to share knowledge when they derive multilingual predictions from subspaces of individual languages, rather than the shared semantic space. Our work also demonstrates that reinforcing processing in the shared space via activation engineering improves multilingual reasoning performance and knowledge transfer in smaller models.
Despite the contributions of this study, we address two limitations. For simplicity, we adopt logit lens to evaluate answer choices in phrases. Other alternatives allow for more expressive interpretations, but require model training [3, 6]. While there is noticeable gain in performance with steered models, the method is only tested on multi-choice questions, and requires extraction of specific activation vectors for individual languages, which may not generalize to other use cases. We also caution against the use of English-centric methods beyond reasoning tasks, which might reinforce cultural and linguistic biases in LLMs [28, 25, 37]. It is however possible for future approaches to learn more stable internal representations when processing multilingual queries, which allow for complete retrieval of knowledge learned in different languages.
Acknowledgments
Special thanks to Grace Chung, Honglin Yu, Alan Ansell and Fan Jiang for their insights and research support.
@@ Appendix B Compute resources
Appendix C Additional results on cross-lingual activation steering
C.1 Hidden activations in Qwen models
@@
Figure 6: Activations in Qwen models on 50 parallel questions from Belebele, analogous to Figure 3 in the main text.
C.2 Varying multiplier on steered Qwen models
@@
C.3 Steering Gemma models across layers
@@
C.4 Accuracy on mHellaSwag, positive transfer and consistency changes in steered models.
Steered results on mHellaSwag is reported in Figure 9.
Figure 9: Accuracy changes in mHellaSwag. The improvement is more substantial in smaller models.
We show the changes in positive transfer from English and consistency with English in Figures 10 and 11 respectively. Overall, the results demonstrate that cross-lingual steering is effective on smaller models; whereas it is more difficult to induce stable and consistent transfer in larger models with simple activation addition.

### v3

@@ 4.2 Multilingual processing in the latent space
(b) Mean accuracy in the latent space across languages for mMMLU. Correct answers can be partially identified in the hidden states. Chance-level accuracy is marked with a dashed line. Full results with the other data sets are reported in Appendix D.
@@
Our next analyses explore multilingual processing with the normalized latent probabilities of answer choices in their native form (nat.) and their English equivalents (en); based on Equation 12.
Figure 2(a) shows the log ratios between $p(\textrm{nat. choices})$ and $p(\textrm{en choices})$, and reveals that larger models tend to assign more mass to answers in the native form in the intermediate layers (ratio above zero). At layer 22 and beyond, the probability of the native language exceed that of English in Gemma 27B. We see a similar trend in Qwen 7B and 32B with English at layers 4 and 40 respectively. This indicates that large-scale language models operate in individual space that is closer to the token embeddings of the native language. The opposite behavior is observed for smaller models: Gemma 2B/9B and Qwen 3B. Early layers of the models are similar to the embeddings of the native languages. The hidden states then shift to working predominantly in the English in the middle layers, where the ratio is below zero, before moving closer to the native language space again at the highest layers. This is in line with previous findings (Tang et al., 2024; Zhao et al., 2024b) that most language-specific neurons are developed at the top and bottom layers of LLMs.
We further interpret the implicit steps taken across layers, by computing for all languages the accuracy of the native and English choices in the latent space. In Figure 2(b) (and Figure 5, Appendix D), the latent predictions are shown to be noisy, however, exhibit accuracy above chance level in higher layers across data sets. Nonetheless, English latent predictions appear more accurate and stable in most cases.
Together with the results in Section 4.1, our analyses imply that smaller models are limited to working within a shared representation, that is closest to English. Conversely, larger models transform the hidden states into subspace of individual languages, rather than the dominant representation. While it is possible for models to reason in a particular language, the process also hinders the models’ ability to infer from the other languages, contributing to cross-lingual inconsistency.
@@ 4.3 Modulating cross-lingual consistency with activation steering
Figure 3: Steered Gemma 2B. Note that $\gamma>0$ indicates an activation vector that pushes the model towards English processing in the latent space.
To establish a causal relation between English-aligned representation and transfer performance, we extract steering vectors from the 50 randomly sampled parallel questions from each data set, and evaluate the cross-lingual performance of the steered models on held-out data. [^5] We reused the training set from Section 4.1.666Prior to steering models, we perform PCA (Wold et al., 1987) to test if the representations are linearly separable, which predetermines the efficacy of the method. We document the results in Appendix E.
Figure 3 reports a sweep of the steering vector multiplier, $\gamma$, on layers 6 and 10 of Gemma 2B. Positively steered models become increasingly accurate (peaking at $2\leq\gamma\leq 3$), are more consistent with English and receive more positive transfer from English. [^7] We observe generally consistent variability across $\gamma$, in terms of standard deviations and confidence intervals over languages. This is omitted in the graph to avoid visual clutter. Similar plots for Qwen models are included in Appendix F.1.
We also observe that steering vectors at the middle layers are the most impactful, because the hidden representations are intended to be aligned at this stage (Figure 6, Appendix E), to allow for language-agnostic task solving (Zhao et al., 2024b). [^8] The performance of steered Gemma models across layers is reported in Figure 9 in Appendix F.2. In Appendix F.3, Figure 10 shows the accuracy changes (%) in steered models across languages. In general, low-resource (accuracy) languages written in Latin script benefit the most from the method in Belebele; whereas the improvement in mMMLU and mHellaSwag is less notable without steering vectors specialized to each topics. The method is also less effective on larger models. It is possible that the hidden representations become increasingly non-linear with scale, leading to more structurally different semantic spaces that are farther apart, as we have shown in Figure 1. Linear perturbations with steering vectors might therefore induce noise and break the geometric structure in the target space of larger models.
@@ 5 Related work
Qi et al. (2023) define pairwise cross-lingual knowledge consistency as the equality in ranked outputs scaled by softmax, whereas Ifergan et al. (2024) and Goldman et al. (2025) measure the extent to which a language pair are accurate in a similar way with factual questions. Our use of Spearman’s correlation in the evaluation of models’ consistency allows a more direct approach, while also taking into account the relative order of ranks. Across languages, script similarity has been proposed to be a major factor to drive knowledge transfer (Qi et al., 2023; Ifergan et al., 2024; Goldman et al., 2025). Our work also supports previous findings that models are capable of processing and storing information with independent representation of different languages (Ifergan et al., 2024), and explains previous findings that transfer performance increases with model size only up to 14B for Qwen 2.5, indicating the ability of larger models to retrieve knowledge learned from independent languages (Goldman et al., 2025).
@@
A growing body of work has shown that multilingual LLMs use English as a pivot language in the hidden layers (Wendler et al., 2024), and share the subspace across languages and modalities (Wu et al., 2024). Alabi et al. (2024) find that the phenomenon also applies to models adapted to new languages, where finetuned adapters preserve most of the representation of the source language of the model, rather than working in a parallel isolated structure. Other works suggest the distinction between language-agnostic and language-specific neurons (Zhao et al., 2024b; Wang et al., 2024), motivating a series of methods that improve multilingual representation and performance of LLMs with targeted supervision (Zhao et al., 2024a; Huo et al., 2025; Wang et al., 2025b). The current study extends this understanding of multilingual models, and provides evidence for larger LLM’s capability to encode meaningful representation in languages beyond English.
@@
Prior work has shown that explicit alignment with English reasoning improves multilingual performance. Existing approaches include English translations of multilingual queries (Shi et al., 2023; Zhu et al., 2024; Ko et al., 2025), using English reasoning traces (Shi et al., 2023; Qin et al., 2023; Yong et al., 2025; Ko et al., 2025), post-hoc translation from English answers (Ko et al., 2025), and trained alignment with English representation (Yoon et al., 2024; Huang et al., 2024; She et al., 2024). Our exploration on cross-lingual activation steering adds to this line of work, and offers a straightforward method to tap into English reasoning capability without training or additional inference costs.
@@ 6 Discussion
This work investigates the challenges of cross-lingual transfer and consistency in LLMs. We examine the implicit processing of multilingual queries and show languages with diverging representations from the shared semantic space exhibit less consistency and knowledge transfer. With more capacity, larger models show a higher utilization of independent subspaces of individual languages, whereas smaller models are constrained to work predominantly in the shared space. Our work also demonstrates that reinforcing processing in the shared space via activation engineering improves multilingual reasoning performance and knowledge transfer in smaller models.
Despite the contributions of this study, we address several limitations. While there is a noticeable gain in performance with steered models, we only apply steering to establish causal evidence for our hypothesis. The method is not meant to be a practical alignment method, as it is only tested on multi-choice questions, and requires extraction of specific activation vectors for individual languages, which may not generalize to other use cases. We also caution against the use of English-centric methods beyond reasoning tasks, which might reinforce cultural and linguistic biases in LLMs (Singh et al., 2024; Rystrøm et al., 2025; Yong et al., 2025). It is however possible for future approaches to learn more stable internal representations when processing multilingual queries, which allow for complete retrieval of knowledge learned in different languages.
Ethics statement
We use steering for the specific purpose of testing the causal relationship proposed by our hypotheses. A significant limitation of this method is its potential to reinforce existing bias towards English in LLMs, which may lead to unfair or harmful outcomes for non-English speakers. We thus recommend future research to address this limitation and to contribute to more equitable multilingual models.
@@ Appendix B Compute resources
Appendix C Impact of language relations on cross-lingual transfer
Figure 4 shows the pairwise positive transfer for all models on mMMLU. Like Ifergan et al. (2024), we find languages with known relations, such as Indic languages: Hindi (hi) and Bengali (bn); and Romance languages: Portuguese (pt), Spanish (es) and Italian (it); to be more closely connected and transfer more knowledge with each other. Indonesian (id) has higher rates of transfer with European languages, relative to Chinese (zh), Korean (ko) and Japanese (ja); indicating the role of shared scripts in knowledge transfer. The models are also less capable of generalizing predictions to and infer knowledge from low-resource languages: Swahili (sw) and Yoruba (yo). Knowledge learned in other languages remains inaccessible in yo as the models improve with increasing capacity.
Appendix D Accuracy in the latent space
Figure 5 shows full results of latent accuracy across languages.
Figure 5: Full results of mean accuracy in the latent space across languages, analogous to Figure 2(b). Chance-level accuracy is marked with a dashed line.
Appendix E PCA visualization of hidden representations
(a) PCA projections of layer activations in Gemma 2 9B.
(b) PCA projections of layer activations in Gemma 2 27B.
Figure 6 provides a visual comparison of the hidden activations of 50 parallel Belebele questions between Gemma 9B and 27B, where the dimension of the vectors is reduced with the projection of Principal Component Analysis (PCA) (Wold et al., 1987). PCA is useful to assess if the hidden representations are linearly separable, which predetermines the efficacy of steering vectors (Rimsky et al., 2024). While the activations can be grouped by languages in the early layers in Gemma 9B (Figure 6(a)), they are mostly aligned between layers 18 and 38, and are separated by their written scripts (Latin vs. non-Latin) in the final layer. We note however that activations of low-resource languages, such as Yoruba (blue squares), are the last to be mapped and continue to be disjoint from the other languages across hidden layers.
@@ Appendix E PCA visualization of hidden representations
Figure 7: Activations in Qwen models on 50 parallel questions from Belebele, analogous to Figure 6.
Appendix F Additional results on cross-lingual activation steering
F.1 Varying multiplier on steered Qwen models
@@
F.2 Steering Gemma models across layers
@@
F.3 Accuracy, positive transfer and consistency changes in steered models across tasks.
(a) Steered models ($\gamma=2.0$) evaluated on Belebele.
(b) Steered models ($\gamma=2.0$) evaluated on mMMLU.
(c) Steered models ($\gamma=2.0$) evaluated on mHellaSwag.
We show the changes in positive transfer from English and consistency with English in Figures 11 and 12 respectively. Overall, the results demonstrate that cross-lingual steering is effective on smaller models; whereas it is more difficult to induce stable and consistent transfer in larger models with simple activation addition.